I am having difficulty setting up this word problem. If a taxi charges .50 for the first 1/5 of a mile, and .20 for every 1/10 of a mile after that, how much would it cost to go 2 miles?
Answer by checkley77(12844)

1/5 mile=.50
2-1/5=1 4/5 mile or 9/5 or 18/10 miles.
18*.20=3.60 for the rest of the 2 mile ride.
Thus the total for the 2 mile ride=.50+3.60=$4.10